OpenStudy (matlee):

Help i need help with yang mills existence and gap , can someone help me solve it

OpenStudy (anonymous):

i can

OpenStudy (matlee):

@satellite73

OpenStudy (matlee):

ty crazy

OpenStudy (anonymous):

ok

OpenStudy (chosenmatt):

I can @matlee

OpenStudy (matlee):

Ok ty chosen

OpenStudy (matlee):

this is what i got so far 3^{3n}+[2(3^{n})]^{3}=3^{3n+2}; n \ge1 (a^{n}-1)^{2n}+(a^{n}-1)^{2n+1}=[a(a^{n}-1)^{2}]^{n}; a \ge2, n \ge3 and [a(a^n+b^n)]^n+[b(a^n+b^n)]^n=(a^n+b^n)^{n+1}; a \ge1, b \ge1, n \ge3 Furthermore, for each solution (with or without coprime bases), there are infinitely many solutions with the same set of exponents and an increasing set of coprime bases. That is, for solution A_1^{x}+B_1^{y}=C_1^{z} we additionally have A_{n}^{x}+B_{n}^{y}=C_{n}^{z}; n \ge2 where A_{n}= (A_{n-1}^{yz+1}) (B_{n-1}^{yz }) (C_{n-1}^{yz }) B_{n}= (A_{n-1}^{xz }) (B_{n-1}^{xz+1}) (C_{n-1}^{xz }) C_{n}= (A_{n-1}^{xy }) (B_{n-1}^{xy }) (C_{n-1}^{xy+1}) Any solutions to the Beal conjecture will necessarily involve three terms all of which are 3-powerful numbers, i.e. numbers where the exponent of every prime factor is at least three. It is known that there are an infinite number of such sums involving coprime 3-powerful numbers;[8] however, such sums are rare. The smallest two examples are: \begin{align} 271^3 + 2^3 3^5 73^3 = 919^3 &= 776,151,559 \\ 3^4 29^3 89^3 + 7^3 11^3 167^3 = 2^7 5^4 353^3 &= 3,518,958,160,000 \\ \end{align}

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

ok.......795415385937

OpenStudy (matlee):

Wow thanks!

OpenStudy (anonymous):

no problem :)